A Bibliography of Publications of Christopher Hugh Bryant

Christopher Hugh Bryant The Robert Gordon University

School of Computing St Andrew St, Aberdeen

AB25 1HG Scotland, UK

Tel: +441224 262737 FAX: +441224 262727

E-mail: chb@scms.rgu.ac.uk (Internet)

25 November 2011 Version 1.05

Abstract

This bibliography records publications of Christopher Hugh Bryant.

Title word cross-reference

10th [CF00]. 11th [LP00].

7th [LD97].

Accuracy [MBS00b]. Active [BMPS99]. Advantage [MBS00b]. Application [BAT^{+}95a]. Available [BAT^{+}95a].

Barcelona [LP00]. Biological [MBS00a].

Case [MBS00b]. Case-study [MBS00b].

Catalonia [LP00]. Chemical [BAT^{+}95b]. Chemistry [BAT^{+}95a]. Chomsky [MBS00a]. Chomsky-like [MBS00a]. Chromatography [BATR94]. close [BMPS99]. Combining [BMPS99, BMO^{+}01]. Commercially [BAT^{+}95a]. Completion [MB00]. Computer [Bry97a]. Conference [CF00]. Czech [LD97].

Data [BAT^{+}95a, BAT^{+}95b]. Database [BAT^{+}95b]. Databases [BR98]. DataMariner [BAT^{+}95a]. Developing [RKK^{+}01]. Discover [BATR97]. Discovering [BAT^{+}95b]. Discovery [BR98]. Domain [BAT^{+}95a].

ECML [LP00].

Enantioseparations [BATR96]. Entailment [MB00]. European [LP00]. Expands [WBM97]. Expert [BATR94].

Families [MBS00a]. Formal [MPWB96]. Function [BMO^{+}01].

General [WBM97]. Generation [Bry97a]. Genes [BMO^{+}01]. Grammars [MBS00a]. grammatical [MBS^{+}01].

Hidden [BAT^{+}95b].

ICML [Lan00]. ICML-2000 [Lan00]. ILP [Bry97b]. ILP-97 [LD97]. Induction [BATR96]. Inductive [BMO^{+}01, BATR97]. international [LD97]. Interpreter [WBM97]. Inverse [MB00].

July [CF00]. June [Lan00].

Knowledge [BAT^{+}95b, BR98].

Learning [BATR96, BMPS99]. like [MBS00a]. Literals [WBM97]. Logic [BATR97, BMO^{+}01]. Logical [RKK^{+}01]. loop [BMPS99].

Machine [BATR96, BMPS99]. May [LP00]. Measuring [MBS00b]. Meta [WBM97]. Metabolism [RKK^{+}01]. Mining [BAT^{+}95a, BAT^{+}95b]. Model [RKK^{+}01].

Negative [WBM97].

Package [BAT^{+}95a]. Performance [MBS00b]. Positives [MBS00b]. Prague [LD97]. Predictive [MBS00b]. proceedings [LD97]. Progol [Bry97b]. Program [WBM97]. Programming [BMO^{+}01, BATR97]. Proofs [WBM97].

Rare [MBS00b]. Relative [MBS00b].

representations [MBS^{+}01].

Republic [LD97].

Requirements [MPWB96].

Review [BATR94]. Robotics [BMO^{+}01]. Rules [BATR96].

September [LD97]. Sequence [MBS00a]. Seventeenth [Lan00]. Spain [LP00]. Specifications [MPWB96]. Stanford [Lan00]. System [BATR96]. Systems [BATR94].

Theory [MB00]. Tool [BAT^{+}95b].

Transforming [WBM97].

UK [CF00]. University [Lan00]. useful [MBS^{+}01]. Using [BAT^{+}95b].

Validation [MPWB96]. versus [MBS00b]. via [Bry97b].

which [WBM97]. workshop [LD97].

Yeast [RKK^{+}01].

References

Bryant:1995:DCA

[BAT^{+}95a] C. H. Bryant, A. E. Adam, D. R. Taylor, G. V. Conroy, and R. C. Rowe. DataMariner, a commercially available data mining package, and its application to a chemistry domain. In Data Mining. UNICOM, London, UK, 1995.

Bryant:1995:DKH

[BAT^{+}95b] C. H. Bryant, A. E. Adam, D. R. Taylor, G. V. Conroy, and R. C. Rowe. Discovering knowledge hidden in a chemical database using a commercially available Data Mining tool. In Knowledge Discovery in Databases, number Digest 1995/021(B) in IEE Computing and Control Division. ????, Savoy Place, London, WC2R OBL, UK, 1995.

Bryant:1994:RES

[BATR94] C. H. Bryant, A. E. Adam, D. R.

Taylor, and R. C. Rowe. Review of expert systems for chromatography. Analytica Chimica Acta, 297(3):317–347, 1994. CO-DEN ACACAM. ISSN 0003-2670.

URL ftp://www.comp.rgu.ac.
uk/pub/staff/chb/bryant_aca_
review.ps.gz.

Bryant:1996:TES

[BATR96] C. H. Bryant, A. E. Adam, D. R.

Taylor, and R. C. Rowe. Towards an expert system for enantioseparations: Induction of rules using machine learning. Chemometrics and Intelligent Laboratory Systems, 34(1):21–40, 1996. CO-DEN CILSEN. ISSN 0169-7439.

URL ftp://www.comp.rgu.ac.
uk/pub/staff/chb/bryant_DataMariner.
ps.gz.

Bryant:1997:UIL

[BATR97] C. H. Bryant, A. E. Adam, D. R.

Taylor, and R. C. Rowe. Using Inductive Logic Programming to discover knowledge hidden in chemical data. Chemometrics and Intelligent Laboratory Systems, 36(2):111–123, 1997. CO-DEN CILSEN. ISSN 0169-7439.

URL ftp://www.comp.rgu.ac.
uk/pub/staff/chb/bryant_golem.
ps.gz.

Bryant:2001:CIL

[BMO^{+}01] C. H. Bryant, S. H. Muggleton, S. G. Oliver, D. B. Kell, P. Reiser, and R. D. King. Combining Inductive Logic programming, Active Learning and robotics to discover the function of genes. Electronic Transactions on Artificial Intelligence, 5(B):1– 36, 2001. CODEN ???? ISSN

                                        ???? URL http://www.ep.liu.
                                        se/ej/etai/2001/001/.

Bryant:1999:CAL

[BMPS99] C. H. Bryant, S. H. Muggleton, C. D. Page, and M. J. E. Sternberg. Combining Active Learning with Inductive Logic Programming to close the loop in Machine Learning. In S. Colton, editor, Proceedings of AISB’99 Symposium on AI and Scientific Creativity, pages 59–64. The Society for the Study of Artificial Intelligence and Simulation of Behaviour (AISB), 1999.

                                        URL ftp://www.comp.rgu.ac.
                                        uk/pub/staff/chb/bryant_aisb99.
                                        ps.gz;   http://www.cogs.susx.
                                        ac.uk/aisb/.

Bryant:1998:KDD

[BR98] C. H. Bryant and R. C. Rowe. Knowledge Discovery in Databases: Application to chromatography. Trends in Analytical Chemistry, 17:18–24, 1 1998. CO-DEN TTAEDJ. ISSN 0165-9936.

                                        URL ftp://www.comp.rgu.ac.
                                        uk/pub/staff/chb/bryant_TRAC.
                                        ps.gz.

Bryant:1997:CGR

[Bry97a] C. H. Bryant. Computer generation of rules for an expert system for enantioseparations. Invited presentation given at Chrial Technology and Enantioseparations ’97, Cambridge, UK, April 1997.

Bryant:1997:DMI

[Bry97b] C. H. Bryant. Data Mining via ILP: The application of Progol to a database of enantioseparations. In Lavrac and Dzeroski [LD97], pages 85–92. CO-DEN LNCSD9. ISBN 3-540-63514-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63.I52 1997.

URL ftp://www.comp.rgu.ac.
uk/pub/staff/chb/bryant_ilp97.
ps.gz.

Cussens:2000:ILP

[CF00] James Cussens and Alan Frisch, editors. Inductive logic programming: 10th International Conference, ILP 2000, London, UK, July 2000: proceedings, volume 1866 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67795-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1866.

Langley:2000:PSI

[Lan00] Pat Langley, editor. Proceedings of the Seventeenth International Conference on Machine Learning (ICML-2000), June 29–July 2, 2000, Stanford University. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2000. ISBN 1-55860-707-2. LCCN Q325.5 .I57 2000.

Lavrac:1997:ILP

[LD97] Nada Lavrac and Saso Dzeroski, editors. Inductive logic programming: 7th international workshop, ILP-97, Prague, Czech Republic, September 17–20, 1997: proceedings, volume 1297 of Lecture Notes in Artificial Intelligence and Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNCSD9. ISBN 3-540-63514-9 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.63.I52 1997.

LopezdeMantaras:2000:MLE

[LP00] Ramon Lopez de Mantaras and Enric Plaza, editors. Machine learning: ECML 2000: 11th European Conference on Machine Learning, Barcelona, Catalonia, Spain, May 31–June 2, 2000, volume 1810 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67602-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1810.

Muggleton:2000:TCU

[MB00] S. H. Muggleton and C. H.

Bryant. Theory completion using inverse entailment. In Cussens and Frisch [CF00], pages 130–146. ISBN 3-540-67795-X (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1866.

URL ftp://www.comp.rgu.ac.
uk/pub/staff/chb/bryant_ilp2k.
ps.gz;   http://www.springer.
de/comp/lncs/index.html.

Muggleton:2000:LCL

[MBS00a] S. H. Muggleton, C. H. Bryant, and A. Srinivasan. Learning Chomsky-like grammars for biological sequence families. In Langley [Lan00], pages 631–638. ISBN 1-55860-707-2. LCCN Q325.5 .I57 2000.

URL ftp://www.comp.rgu.ac.
uk/pub/staff/chb/bryant_icml2k.
ps.gz.

Muggleton:2000:MPW

[MBS00b] S. H. Muggleton, C. H. Bryant, and A. Srinivasan. Measuring performance when positives are rare: Relative advantage versus predictive accuracy — a biological case-study. In Lopez de Mantaras and Plaza [LP00], pages 300–312. ISBN 3-540-67602-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1810.

URL ftp://www.comp.rgu.ac.
uk/pub/staff/chb/bryant_ecml2k.
ps.gz;   http://www.springer.
de/comp/lncs/index.html.

Muggleton:2001:GRU

[MBS^{+}01] S. H. Muggleton, C. H. Bryant, A. Srinivasan, A. Whittaker, S. Topp, and C. Rawlings. Are grammatical representations useful for learning from biological sequence data? — a case study. Journal of Computational Biology, 8(5):493–522, October 2001. CODEN JCOBEM. ISSN 1066-5277. URL http://www.

                                        liebertpub.com/. \begin{matrix}c \\ ◯ \end{matrix}Mary Ann

Liebert.

McCluskey:1996:VFS

[MPWB96] T. L. McCluskey, J. M. Porteous, M. M. West, and C. H. Bryant. The validation of formal specifications of requirements. In Proceedings of the BCS-FACS Northern Formal Methods Workshop, Ilkley, UK, Electronic Workshops in Computing Series. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., September 1996.

                                        URL ftp://www.comp.rgu.ac.
                                        uk/pub/staff/chb/bryant_north_
                                        fm_ws.ps.gz.

Reiser:2001:DLM

[RKK^{+}01] P. Reiser, R. D. King, D. B. Kell, S. H. Muggleton, C. H. Bryant, and S. G. Oliver. Developing a logical model of yeast metabolism. Electronic Transactions on Artificial Intelligence, 5 (B):223–244, 2001. CODEN ????

                                        ISSN ???? URL http://www.ep.
                                        liu.se/ej/etai/2001/013/.

West:1997:TGP

[WBM97] M. M. West, C. H. Bryant, and T. L. McCluskey. Transforming general program proofs: A meta interpreter which expands negative literals. In The preliminary Proceedings of the Seventh International Workshop on Logic Program Synthesis and Transformation. ????, Leuven, Belgium, 1997.

URL ftp://www.comp.rgu.ac.
uk/pub/staff/chb/bryant_lopstr97.
ps.gz.